# Black Hole Complementarity and ER/EPR

### v1

@@ Abstract
Essay written for the Gravity Research Foundation 2025 Awards for Essays on Gravitation.
@@ 1 Introduction
It is this problem that we will address in the present work. In particular, we will make use of tools from both relativity and quantum information—specifically, traversable wormhole constructions and quantum channels—to show that all wormholes in a common exterior spacetime must share entanglement between their exterior degrees of freedom, regardless of asymptotic boundary conditions, provided that black hole complementarity is correct [26].
@@ 2.2 Review of Black Hole Complementarity
Complementarity has come under criticism over the past fifteen years in the context of the firewall paradox [50, 51, 52], a sharpening of the black hole information puzzle. We will not review the information paradox here, save to say that the currently favored resolution via entanglement islands [53, 54] alleviates much of the direct tension with black hole complementarity.
@@ 3 Complementarity Implies Entangled Wormholes
We can lower-bound the necessary entanglement present in the wormhole as follows. The one-qubit channel capacity $C_{1}(\varepsilon)$—that is, the maximal amount of information in qubits conveyed by the channel $\varepsilon$ when one qubit is sent through from a system $L$ to $R$—is upper-bounded by $1$. More broadly, in entanglement-assisted quantum channels, the channel capacity for an $n$-qubit transmission is upper-bounded by $n$, which is by definition upper-bounded by the number of Bell pairs $N$ distillable between $L$ and $R$. Finally, writing the von Neumann entropy $S(A)=-{\rm Tr}\rho_{A}\log\rho_{A}$, we have that the mutual information between subsystems,
$I(L:R)=S(L)+S(R)-S(LR),$ (7)
@@
so since our traversable wormhole achieves some nonzero transmission of information via our signal ${\cal T}$, from the black hole complementarity perspective of the exterior observer, there must be nonzero mutual information between the two mouths $L$ and $R$. We emphasize that this conclusion is reached without appeal to AdS/CFT, but directly as a consequence of black hole complementarity in a spacetime with arbitrary asymptotics. While it is possible to consider quantum channels associating more than two subsystems, for this work it is sufficient to restrict to two subsytems, and thus the number of distillable Bell pairs is sufficient to characterize the entanglement, rather than more general forms of squashed entanglement.
@@ 4 Entangled Horizons as an Initial Condition
We have shown that, once rendered traversable, wormholes must have stretched horizons that are entangled with each other. We will now argue that this entanglement must be intrinsic to the wormhole itself—rather a byproduct of traversability—so that the entanglement predated the null shock wave. That is, all two-sided wormholes, traversable or not, must be entangled if the two sides are in the same asymptotic spacetime.
@@
For the wormholes we consider, the two stretched horizons must necessarily be in causal contact, in order for us to be permitted to make the wormhole traversable without violating the achronal ANEC as discussed in Sec. 2.1. While causally connected regions of space in quantum field theory are always entangled in the vacuum, by virtue of the Reeh-Schlieder theorem [58], this conclusion does not apply to entanglement of the wormhole mouths for two reasons. First, the state in the exterior need not necessarily be the vacuum; further, and more crucially, the Reeh-Schlieder theorem applies to local relativistic quantum field theories, i.e., to Type III von Neumann algebras. We note that the von Neumann algebra relevant to stretched horizons of black holes is Type I, as the number of degrees of freedom is finite due to the finiteness of Bekenstein-Hawking entropy [59, 60]. A modern take on this problem by Witten et al. [61] in the context of the timelike tube theorem similarly shows that the von Neumann algebra type changes once coupled to gravity [62]. Even if some weaker version of Reeh-Schlieder applied morally in the wormhole setup, one still could not prove that the entanglement implied would be usable for any quantum information theoretic task; indeed, it may be analogous to the unusable entanglement of the fermionic vacuum [63].
@@
$||[{\cal O}_{L}(t),{\cal O}_{R}]||\leq c\,e^{-\lambda(d(L,R)-v_{LR}|t|)},$ (9)
where $d(L,R)$ is the proper distance and |||||$\cdot$||| | ⋅ | | is the operator norm [64]. Estimates of the butterfly velocity $v_{B}$—essentially an effective Lieb-Robinson velocity for chaotic systems—have been made [65, 66], and in large-$N$ holographic systems one has $v_{B}=\sqrt{\frac{D+1}{2D}}$. While we will not use the precise holographic result, as long as $v_{LR}<1$, quantum gravitational back reaction is precluded from being able to generate the entanglement at the requisite separation. The only remaining option is that the stretched horizons of the wormhole were entangled before the traversability process began, simply by virtue of having a shared interior.
@@ Acknowledgments
We thank Aidan Chatwin-Davies and Aidan Herderschee for useful comments. N.B. is supported by Northeastern University Department of Physics and by Brookhaven National Laboratory. G.N.R. is supported by the James Arthur Postdoctoral Fellowship at New York University.

### v2

@@ Abstract
Honorable Mention, Gravity Research Foundation 2025 Awards for Essays on Gravitation
@@ 1 Introduction
It is this problem that we will address in the present work. In particular, we will make use of tools from both relativity and quantum information—specifically, traversable wormhole constructions and quantum channels—to show that all wormholes in a common exterior spacetime must share entanglement between their exterior degrees of freedom, regardless of asymptotic boundary conditions, provided that black hole complementarity is correct [26]. While the true entanglement structure of horizons is itself an extremely difficult problem in quantum gravity, as sharpened by the firewall formulation of the black hole information paradox [27, 28, 29], it will be striking that, crucially provided that complementarity holds in a sense that we will make clear, the general conclusion that horizons with a shared interior must be entangled follows from a weaker set of assumptions than AdS/CFT alone.
@@ 2.2 Review of Black Hole Complementarity
Complementarity has come under criticism over the past fifteen years in the context of the firewall paradox [27, 28, 29]. While some in the community have proposed that entanglement islands [53, 54] can resolve aspects of the information paradox in a way that has been suggested to preserve black hole complementarity, consensus has not yet been reached [55, 56, 57, 58, 59, 60, 61] as to whether islands fully solve the information paradox in its modern firewalls form or whether complementarity itself fully survives.
@@ 3 Complementarity Implies Entangled Wormholes
The fact that LOCC between observers outside of a traversable wormhole is equivalent to a quantum channel in the complementary frame was established by Gao, Jafferis, and Wall in the context of AdS/CFT [33]. In our context, we will not be using AdS/CFT but instead the broader framework of black hole complementarity. As we have emphasized, in black hole complementarity all processes behind the horizon correspond to some process on the quantum state defined on the stretched horizons. The most general process on any quantum state is a quantum channel, so without loss of generality, we can view the transmission of information through a traversable wormhole, in the black hole complementarity picture, as some quantum channel acting on the two stretched horizons. As should be clear, this interpretation hinges crucially on the validity of black hole complementarity, so if the final resolution of the information paradox turns out to violate complementarity, our conclusions will not hold. In particular, the consumption of entanglement in an entanglement assisted quantum channel always relies on a measurement occurring [63]. While the explicit form of such a measurement in the stretched horizon picture is not immediately obvious, in the complementary frame of the black hole interior, the negative null energy shock wave plays this role by definition—as explicitly explored in Ref. [33]—since the shock wave is the only object in the interior that is not the message itself. We leave a detailed exploration of the measurement problem and decoherence on the stretched horizon side of the duality in black hole complentarity to future work.
@@
We can lower-bound the necessary entanglement present in the wormhole as follows. The one-qubit channel capacity $C_{1}(\varepsilon)$—that is, the maximal amount of information in qubits conveyed by the channel $\varepsilon$ when one qubit is sent through from a system $L$ to RR—is upper-bounded by $1$. More broadly, in entanglement-assisted quantum channels, the channel capacity for an $n$-qubit transmission is upper-bounded by $n$, which is by definition upper-bounded by the number of Bell pairs $N$ distillable between $L$ and $R$. Finally, writing the von Neumann entropy $S(A)=-{\rm Tr}\rho_{A}\log\rho_{A}$, we have that the mutual information between subsystems,
I(L:R)=S(L)+S(R)S(LR),I(L:R)=S(L)+S(R)-S(LR), (7)
@@
so since our traversable wormhole achieves some nonzero transmission of information via our signal $\cal T$}, from the black hole complementarity perspective of the exterior observer, there must be nonzero mutual information between the two mouths $L$ and RR. [^2] In particular, we expect that the maximal amount of information transmitted—and hence our expected lower bound on the entanglement between the horizons—scales like some small parameter $\delta\ll 1$ times ${\cal A}/4G$ for $\cal A$} of order the area of the horizons, since we can transmit $\delta\times O({\cal A}/4G)$ qubits while still remaining in perturbative control of the geometry. We leave further quantification of the horizon entanglement to future work. We emphasize that this conclusion is reached without appeal to AdS/CFT, but directly as a consequence of black hole complementarity in a spacetime with arbitrary asymptotics. While it is possible to consider quantum channels associating more than two subsystems, for this work it is sufficient to restrict to two subsytems, and thus the number of distillable Bell pairs is sufficient to characterize the entanglement, rather than more general forms of squashed entanglement.
@@ 4 Entangled Horizons as an Initial Condition
We have shown that, once rendered traversable, [^3] Recall that in Ref. [43], the protocol for generating the NEC-violating shock wave that makes the wormhole traversable is through Casimir effects when the two horizons are sufficiently close together; one generates a NEC-violating shock wave by simply taking two horizons with a shared interior and gravitationally towing them to be close together in the exterior spacetime. In Ref. [33], the shock wave protocol is instead the turning on of the double trace deformation coupling in the AdS/CFT dual. wormholes must have stretched horizons that are entangled with each other. We will now argue that this entanglement must be intrinsic to the wormhole itself—rather a byproduct of traversability—so that the entanglement predated the null shock wave. That is, all two-sided wormholes, traversable or not, must be entangled if the two sides are in the same asymptotic spacetime.
@@
For the wormholes we consider, the two stretched horizons must necessarily be in causal contact, in order for us to be permitted to make the wormhole traversable without violating the achronal ANEC as discussed in Sec. 2.1. While causally connected regions of space in quantum field theory are always entangled in the vacuum, by virtue of the Reeh–Schlieder theorem [65], this conclusion does not apply to entanglement of the wormhole mouths for two reasons. First, the state in the exterior need not necessarily be the vacuum; further, and more crucially, the Reeh–Schlieder theorem applies to local relativistic quantum field theories, i.e., to Type III von Neumann algebras. We note that the von Neumann algebra relevant to stretched horizons of black holes is Type I, as the number of degrees of freedom is finite due to the finiteness of Bekenstein-Hawking entropy [66, 67]. A modern take on this problem by Witten et al. [68] in the context of the timelike tube theorem similarly shows that the von Neumann algebra type changes once coupled to gravity [69]. Even if some weaker version of Reeh–Schlieder applied morally in the wormhole setup, one still could not prove that the entanglement implied would be usable for any quantum information theoretic task; indeed, it may be analogous to the unusable entanglement of the fermionic vacuum [70]. Indeed, any argument that would imply that Reeh–Schlieder-like entanglement in the ambient spacetime surrounding each observer is responsible for the entanglement assisted quantum channel that enables their communication would equally well apply to two horizons that do not have a shared interior, and communication between such horizons—in the absence of LOCC through the exterior spacetime, which we have forbidden—is not possible. Moreover, we can always idealize our observer as a finite-dimensional quantum system, simply by considering a quantum computer with a finite number of qubits coupled to the stretched horizon. While we could consider a field-theoretic observer endowed with a Type III algebra, it suffices to demonstrate the necessity of entanglement between the two horizons in the case of finite-dimensional observers.
@@
${\cal O}_{L}(t),{\cal O}_{R}]||\leq c\,e^{-\lambda(d(L,R)-v_{LR}|t|)},$ (9)
where $d(L,R)$ is the proper distance and $||\cdot||$ is the operator norm [71]. Estimates of the butterfly velocity $v_{B}$—essentially an effective Lieb-Robinson velocity for chaotic systems—have been made [72, 73], and in large-$N$ holographic systems one has $v_{B}=\sqrt{\frac{D+1}{2D}}$. While we will not use the precise holographic result, as long as $v_{LR}<1$, quantum gravitational back reaction is precluded from being able to generate the entanglement at the requisite separation. The only remaining option is that the stretched horizons of the wormhole were entangled before the traversability process began, simply by virtue of having a shared interior.
@@ Acknowledgments
We thank Aidan Chatwin-Davies, Ahmed Almheiri, and Aidan Herderschee for useful conversations. N.B. is supported by Northeastern University Department of Physics and by Brookhaven National Laboratory. G.N.R. is supported by the James Arthur Postdoctoral Fellowship at New York University.